OpenStudy (cookiemonster18):

8 Geometry questions, Please help me correct my incorrect answers.

OpenStudy (cookiemonster18):

1. Which construction is illustrated above? A. a parallel line to a given line from a point not on the line.***** B. an angle congruent to a given angle. C. the bisector of an angle. D a perpendicular to a given line from a point on the line.

OpenStudy (cookiemonster18):

2. Find the value of x and y. A. x=15, y=12***** B. x=14, y=11 C. x=14, y=12 D. x=15, y=11

OpenStudy (cookiemonster18):

3. What is the relationship between <2 and <7? A. corresponding angles B. same-side interior angles C. alternate exterior angles***** D. alternate interior angles

OpenStudy (cookiemonster18):

4. Construct the line perpendicular to TU at point V. I think its the third picture (picture 2 of 4)*****

OpenStudy (cookiemonster18):

5. For which value of x is line L parallel to line M? A. 28 B. 62***** C. 22 D. 68

OpenStudy (cookiemonster18):

6. Which of the following facts, would allow you to prove that line L and M are parallel? A. m<2=m<3 B. m<7+m<8=180 C. m<4=m<5***** D. m<6+m<2=180

OpenStudy (cookiemonster18):

7. A skateboard ramp has a 158 entry angle. Find the measure of <1 A. 22 B. 68***** C. 78 D. 58

OpenStudy (cookiemonster18):

8. Which pair of angles are alternate exterior angles? A. <1 and <8 B. <1 and <7 C. <2 and <7***** D. <2 and <9

OpenStudy (cookiemonster18):

My answers are: 1. A 2. A 3. C 4. B 5. B 6. C 7. B 8. C (Please help correct me if I'm wrong.)

OpenStudy (seratul):

1 should be B. The angles are congruent 5 should be C. 130+2x+6=180 136+2x=180 2x=44 x=22 Everything else seems right :D Good job.

OpenStudy (cookiemonster18):

For anyone who many need future help in geometry Lesson 9 Unit 4: Parallel and Perpendicular Lines. Here are all the correct answers from 1-8. 1. A 2. A 3. C 4. B 5. C 6. C 7. B 8. C 8/8 100%